OpenStudy (anonymous):

Use a calculator to help solve the problem. (Round your answer to one decimal place.) The concentration x of a certain drug in an organ after t minutes is given by the equation below. Find the concentration of the drug in 3/4 hour. x = 0.09(1 − e−0.2t)

OpenStudy (anonymous):

just put t=3/4 and use calculator to find x...

OpenStudy (anonymous):

you mean t = 3/4 * 60 t = 45

OpenStudy (anonymous):

I'm assuming this is the correct equation \[x(t) = 0.09(1-e^{.02t})\] \[x (45\min) = 0.09*(1-e^{.02*45}) = 0.05\]

OpenStudy (anonymous):

t and x do not multiply on the left side of the =. and it is -.2 not .02

OpenStudy (anonymous):

On the left side, you are incorrect, that doesn't mean multiplication, but you are correct about the .2 vs .02 \[x(t)=0.09(1−e^{.2t})\] \[x(45\min)=0.09∗(1−e^{.2∗45})=0.09\]